Inactive Tutor answered 11/21/20
P(H/R) = 50% If it's red, there are only 2 alternatives, a heart or a diamond
P(H/F) = 25% It's one of 4 suits. 1/4 =25%
P(F/H) = 23.09% 3 face cards out of 13 possible cards = 3/13 =23.09% approximately
Fav D.
asked 11/21/20b. What is the probability of drawing a heart if you know that you have drawn a
red card?
c. What is the probability of drawing a heart if you know that you have a face
card?
d. What is the probability of drawing a face card if you know that you have a
heart?
Inactive Tutor answered 11/21/20
P(H/R) = 50% If it's red, there are only 2 alternatives, a heart or a diamond
P(H/F) = 25% It's one of 4 suits. 1/4 =25%
P(F/H) = 23.09% 3 face cards out of 13 possible cards = 3/13 =23.09% approximately
Get a free answer to a quick problem.
Most questions answered within 4 hours.
Choose an expert and meet online. No packages or subscriptions, pay only for the time you need.